I would appreciate some help on a math problem that has me completely flustered. The question is; The product of two numbers is 60. If one of the numbers is 15/19(fraction), what is the other number?
Answer by Alan3354(69443)

I would appreciate some help on a math problem that has me completely flustered. The question is; The product of two numbers is 60. If one of the numbers is 15/19(fraction), what is the other number?
------------
n = the other number
n*(15/19) = 60
Divide by 15
n*(1/19) = 60/15 = 4
Multiply both sides by 19
n = 4*19
n = 76